# It’s Giving Major Satisfaction: Why Fairness Matters for Developers

### v1

@@ VII-A On Overall Satisfaction and Job Security
VII-B Interpersonal Fairness at the Core
Among all the dimensions influencing overall job satisfaction, interpersonal fairness has the greatest effect. In particular, developers who feel respected are more than twice as likely to report high job satisfaction. Following closely were treatment with dignity, polite treatment, and avoidance of inappropriate comments or behavior. This suggests that out of 4 fairness dimensions influencing overall job satisfaction, interpersonal fairness is at the core.
Fair treatment fosters an environment where individuals can thrive, and this environment is often characterized by high levels of psychological safety [95]. According to Google’s Project Aristotle, which analyzed data from 180 software teams and 35 statistical models, psychological safety is the most important factor for team success [96]. Software teams that exhibit high psychological safety are not only more innovative and effective but also experience lower turnover rates [96].
In the fast-paced and collaborative world of SE, psychological safety allows team members to share knowledge, admit mistakes, and propose innovative solutions without fear of negative consequences [96]. This open and supportive atmosphere is closely linked to how individuals are treated within their teams. This highlights the critical role that interpersonal fairness plays in fostering psychological safety. Our findings are also consistent with prior research. Westlund [97] found that leadership styles incorporating respect and recognition strongly correlate with increased job satisfaction among software developers.
@@ VII-C Acknowledging Contributions
The context of authorship seems particularly well-suited for applying fairness principles, as fairness in this context significantly influences overall satisfaction, rewards, work-life balance, job security, manager, and team culture satisfaction. In software development, it has been studied that not all contributions are equally recognized. While adding new features is often celebrated, critical but less visible tasks, like bug fixing and code reviews, frequently go unacknowledged [1]. This issue is particularly prevalent in open-source ecosystems, where credit attribution remains a challenge.
@@
Peripheral developers, those who engage in tasks like bug fixing, are often underappreciated in traditional attribution systems, which tend to favor core contributors [1]. This uneven distribution of credit reinforces power dynamics and can lead to dissatisfaction among those who contribute but remain unrecognized [1]. Geiger et al. further highlighted that maintainers of open source projects perform significant invisible labor, which is critical for sustaining projects but often goes unrecognized [98]. Moreover, Kononenko et al. have shown that code reviews are sometimes underappreciated [99]. Despite their importance in maintaining software quality, the contributions of those who conduct thorough code reviews can go unnoticed.
@@
To address these imbalances, some open-source ecosystems have begun adopting more equitable models for recognizing contributions. For instance, the All Contributors model seeks to fairly acknowledge work beyond just code, including bug reporting, documentation, and community management [100]. This ensures that contributors beyond the core developers are properly recognized for their efforts. At the same time, several studies have explored methods for identifying authorship in source code for both legal purposes and ensuring fair credit for contributions. Bogomolov et al. introduced a language-agnostic method for code authorship attribution to detect plagiarism and resolve legal disputes [101]. A more recent study by Czibula et al. presented an ensemble of deep autoencoders to improve software authorship attribution to enhance the accuracy in identifying developers behind a code [102].
@@
Adding to the existing literature, this study underscores the importance of applying the distributive, procedural, interpersonal, and informational fairness when recognizing contributions in SE. For example, German et al. applied Colquitt’s fairness theory to create a framework for assessing fairness in modern code review processes [5, 103, 6]. This framework provides actionable guidance on implementing fairness principles across the four dimensions in code reviews.
@@
In terms of distributive fairness, organizations should determine how to measure contributions and decide whether individuals who contribute more should receive greater recognition or rewards [6]. This is to ensure that compensation reflects the value and effort of contributions. Procedural fairness involves establishing clear standards for successful contributions and maintaining transparent, accessible review processes [6]. This includes, for example, allowing contributors to comment on reviews and providing an appeal mechanism for rejected work [6].
@@
Interpersonal fairness is key to maintaining respectful interactions during reviews [6]. By defining and enforcing a code of conduct, organizations can ensure contributors are treated respectfully, addressing any inappropriate behavior quickly [6]. Informational fairness, meanwhile, requires ongoing transparency in the review process. This can involve, for instance, keeping all parties informed, offering constructive feedback, and clearly communicating delays or unexpected outcomes [6]. By embedding these principles into code reviews and other recognition processes, software teams can create a more fair system for acknowledging contributions.
@@
On the other hand, researchers can investigate the impact of various contribution attribution model on team satisfaction, project success, and contributor retention [104, 105]. By analyzing real-world implementations of fairness frameworks, researchers could identify which approaches yield the most positive outcomes and how to adapt them to different team dynamics.
VII-D Satisfaction with Managers
@@
Previous studies support these findings. Suh & Hijal-Moghrabi demonstrated that fair HR practices and fairness from superiors strongly impacted women’s job satisfaction [106]. Choi & Rainey also found that diversity management and organizational fairness positively influenced job satisfaction, particularly for women, who reported higher satisfaction when fairness was perceived [107]. Choi further discovered that racial/ethnic minorities experienced higher job satisfaction when they held majority status in their organization, while Whites reported greater satisfaction in predominantly White settings [108]. This highlights how ethnicity representation can impact one’s job satisfaction.
Gender and ethnic biases are prevalent SE as well. Gunawardena et al. [109] showed that destructive criticism during code reviews negatively affects female developers’ moods and willingness to collaborate. Trinkenreich et al. [110] found that women are sometimes assigned to projects based on gender rather than ability, while Paul et al. [111] revealed that male developers are more likely to write negative comments and offer less assistance to female colleagues. Our study builds on this body of work, which suggests that managers can play a crucial role in addressing gender bias. By prioritizing fair outcomes, clear communication, and unbiased decision-making processes, managers can significantly improve female developers’ satisfaction.
Similarly, developers from underrepresented ethnic groups, who are often a minority in their teams, face additional challenges in feeling heard and valued. Our findings show that for these developers, satisfaction with managers is strongly influenced by how inclusive and transparent managerial practices are, particularly in decision-making and providing opportunities to voice concerns or appeal decisions. When underrepresented developers feel they lack equal opportunities to participate, they may perceive this as unfair treatment, leading to dissatisfaction with their managers. This sense of disadvantage may stem from a lack of inclusion and transparency [112], which can cause these developers to feel excluded from important team decisions.
This aligns with research by Hawes [113], which found that minority workers’ job satisfaction improves significantly with strong supervisory support. In the predominantly White-dominated field of SE, Nadri et al. [3] reported that less than 10% of contributions to open-source projects come from visibly non-White developers. These systemic barriers further emphasize the need for inclusive and transparent managerial practices to ensure underrepresented developers feel empowered and valued in their roles.
@@
VII-E Policy Barriers for Women and Minorities
@@
Policies that lack transparency or rely on subjective criteria can disproportionately disadvantage female and underrepresented developers. These groups often lack access to informal networks that help advance careers, facing a glass ceiling despite their qualifications [114]. Even when women and minorities meet formal criteria like performance ratings, they often struggle to advance because they are excluded from the informal mentorship or sponsorship opportunities that typically benefit others, particularly white males [114]. Unconscious biases continue to hinder their progress in Science, Technology, Engineering, and Mathematics (STEM) fields, limiting their opportunities for promotion and career growth, even in workplaces that have diversity policies in place [115, 116].
Inflexible policies around working hours, parental leave, or caregiving responsibilities can also disproportionately impact female developers. Research from sectors like healthcare shows that work-life balance strongly affects organizational commitment and job satisfaction for women [117]. Similarly, remote work policies pose challenges if they are not equally accessible or if there is stigma attached to using them. Women, particularly those using flexible work options, often face a flexibility stigma, which can lead to perceptions of being less committed [118, 119, 120]. The COVID-19 pandemic has reduced some of the stigma around remote work, offering an opportunity to create more inclusive environments for women and minority employees [121].
@@
VII-F Managing Work Hours for the Less Experienced
@@
Several factors may contribute to the challenges faced by less experienced developers. One key challenge is the experience gap. Junior employees often struggle with unfamiliarity with company processes, technical skills, and organizational culture, which delays their performance and puts them at a disadvantage compared to senior employees who have already navigated these hurdles [87]. Experienced developers tend to introduce fewer defects, even when working on similar tasks, and are generally more efficient in their performance [122]. Their experience and knowledge of the project help them handle complex tasks more efficiently, often completing them faster and with better results [123]. In contrast, less experienced developers spend more time searching for information and understanding code, leading to slower work processes [88]. As less experienced developers gain more experience, their performance gradually become more like those of experienced developers [124].
Research shows that long and irregular working hours negatively impact mental capacity and work quality for all developers, regardless of experience. Extended periods of continuous work reduce efficiency [125]. Meyer et al. suggest developing better tools to help developers plan their workday and boost personal productivity [126]. Additionally, Rastogi et al. highlight the value of mentorship or buddy systems for new hires, which can help them quickly understand company processes and become productive faster [87]. A supportive environment where new employees feel comfortable asking questions without fear of judgment is also important for their development [87].
@@
VII-G Supporting Work-Limited Developers
Our analysis shows that developers with work limitations are often disadvantaged, especially when it comes to compensation in high-pressure situations. According to the National Health Interview Survey (NHIS), a work limitation is defined as a condition that limits a person’s ability to perform certain tasks or activities required for their job [61]. For software engineers and developers, this could mean physical, cognitive, or psychological challenges that may affect their performance, such as chronic illnesses, disabilities, or other long-term health conditions.
A study examining work limitations due to chronic health conditions found that such limitations often led to reduced productivity and lower wages [127]. However, workers whose jobs were better matched to their limitations experienced better outcomes, including higher wages and longer job tenure [127]. Similarly, Siu et al. found that employees with chronic diseases valued fair treatment, clear communication, and flexible accommodations, all of which positively impacted their job satisfaction and work performance [128].
Providing accommodations and adjusting work conditions for employees with chronic physical conditions has been shown to improve their self-efficacy and reduce fatigue [129]. Lerner et al. further confirmed that employees with chronic health conditions face challenges in managing time, physical demands, and interpersonal tasks, often resulting in productivity loss [130]. This aligns with the importance of non-monetary compensation, such as flexible work hours and support services, as essential factors for job satisfaction among these employees [130].
Developers with disabilities face additional challenges in the workplace, including issues with project management practices, workplace dynamics, and the accessibility of development tools. For example, visually impaired developers often struggle with the accessibility of development environments and tools, which negatively impacts their productivity [131, 132]. Programs like Catalisa, which train and hire people with disabilities as software developers, underscore the importance of tailored training to improve inclusion and accessibility within the industry [133].
@@ VII-G Supporting Work-Limited Developers
Future research could explore the types of accommodations that are most effective in improving job satisfaction for software developers with work limitations. This could include case studies of specific organizations that implement flexible working conditions [134, 135, 136], and how different types of accommodations [137] (e.g., flexible hours, ergonomic tools, or assistive technology) impact the developers.
VIII Threats to Validity
Construct Validity. Since perceptions of fairness and job satisfaction are subjective, individual interpretations of survey questions may vary based on personal experiences or biases. To minimize this, we used well-established frameworks, including Colquitt’s fairness theory [5] and Storey et al.’s job satisfaction model [12]. These scales have been rigorously tested and help ensure that the questions measure what they are intended to. We also ensured anonymity in the survey and conducted a pilot test with six respondents to refine any unclear questions and adjust wording based on feedback.
@@ VIII Threats to Validity
Internal Validity. While our study highlights a relationship between fairness perceptions and job satisfaction, we acknowledge that it does not conclusively establish causality. However, this limitation is mitigated by both our calculation of the effect size and the extensive body of organizational justice literature, consistent with prior research across various industries and contexts [30, 31, 32, 8, 9, 10, 7]. Given this robust theoretical grounding and the significant effect sizes of minimum 50% effect observed in our study, we argue that the direction of the relationship identified is aligned with well-established patterns in organizational behavior.

### v2

@@ 7.1. On Overall Satisfaction and Job Security
7.2. Interpersonal Fairness at the Core
Among all the dimensions influencing overall job satisfaction, we found that interpersonal fairness has the greatest effect. In particular, developers who feel respected are more than twice as likely to report high job satisfaction. Following closely were treatment with dignity, polite treatment, and avoidance of inappropriate comments or behavior. This suggests that out of four fairness dimensions influencing overall job satisfaction, interpersonal fairness is at the core.
Fair treatment fosters an environment where individuals can thrive, and this environment is often characterized by high levels of psychological safety (Newman et al., 2017). According to Google’s Project Aristotle, which analyzed data from 180 software teams and 35 statistical models, psychological safety is the most important factor for team success (Osmani, 2024). Software teams that exhibit high psychological safety are not only more innovative and effective but also experience lower turnover rates (Osmani, 2024).
In the fast-paced and collaborative world of SE, psychological safety allows team members to share knowledge, admit mistakes, and propose innovative solutions without fear of negative consequences (Osmani, 2024). This open and supportive atmosphere is closely linked to how individuals are treated within their teams. This highlights the critical role that interpersonal fairness plays in fostering psychological safety. Our findings are also consistent with prior research. Westlund (Westlund, 2011) found that leadership styles incorporating respect and recognition strongly correlate with increased job satisfaction among software developers.
@@ 7.3. Acknowledging Contributions
The context of authorship seems particularly well-suited for applying fairness principles, as fairness in this context significantly influences overall satisfaction, rewards, work-life balance, job security, manager, and team culture satisfaction. In software development, it has been studied that not all contributions are equally recognized. While adding new features is often celebrated, critical but less visible tasks, like bug fixing and code reviews, frequently go unacknowledged (Casari et al., 2021). This issue is particularly prevalent in open-source ecosystems, where credit attribution remains a challenge.
@@
Peripheral developers, those who engage in tasks like bug fixing, are often underappreciated in traditional attribution systems, which tend to favor core contributors (Casari et al., 2021). This uneven distribution of credit reinforces power dynamics and can lead to dissatisfaction among those who contribute but remain unrecognized (Casari et al., 2021). Geiger et al. (Geiger et al., 2021) further highlighted that maintainers of open source projects perform significant invisible labor, which is critical for sustaining projects but often goes unrecognized. Moreover, Kononenko et al. (Kononenko et al., 2015) have shown that code reviews are sometimes underappreciated. Despite their importance in maintaining software quality, the contributions of those who conduct thorough code reviews can go unnoticed.
@@
To address these imbalances, some open-source ecosystems have begun adopting more equitable models for recognizing contributions. For instance, the All Contributors model seeks to fairly acknowledge work beyond just code, including bug reporting, documentation, and community management (Young et al., 2021). This ensures that contributors beyond the core developers are properly recognized for their efforts. At the same time, several studies have explored methods for identifying authorship in source code for both legal purposes and ensuring fair credit for contributions. Bogomolov et al. (Bogomolov et al., 2020) introduced a language-agnostic method for code authorship attribution to detect plagiarism and resolve legal disputes. A more recent study by Czibula et al.(Czibula et al., 2022) presented an ensemble of deep autoencoders to improve software authorship attribution to enhance the accuracy in identifying developers behind a code.
@@
Adding to the existing literature, this study underscores the importance of applying the distributive, procedural, interpersonal, and informational fairness when recognizing contributions in SE. For example, German et al. (German et al., 2018) applied Colquitt’s fairness theory (Colquitt, 2001; Colquitt and Chertkoff, 2002) to create a framework for assessing fairness in modern code review processes. This framework provides actionable guidance on implementing fairness principles across the four dimensions in code reviews.
@@
In terms of distributive fairness, organizations should determine how to measure contributions and decide whether individuals who contribute more should receive greater recognition or rewards (German et al., 2018). This is to ensure that compensation reflects the value and effort of contributions. Procedural fairness involves establishing clear standards for successful contributions and maintaining transparent, accessible review processes (German et al., 2018). This includes, for example, allowing contributors to comment on reviews and providing an appeal mechanism for rejected work (German et al., 2018).
@@
Interpersonal fairness is key to maintaining respectful interactions during reviews (German et al., 2018). By defining and enforcing a code of conduct, organizations can ensure contributors are treated respectfully, addressing any inappropriate behavior quickly (German et al., 2018). Informational fairness, meanwhile, requires ongoing transparency in the review process. This can involve, for instance, keeping all parties informed, offering constructive feedback, and clearly communicating delays or unexpected outcomes (German et al., 2018). By embedding these principles into code reviews and other recognition processes, software teams can create a more fair system for acknowledging contributions.
@@
On the other hand, researchers can investigate the impact of various contribution attribution model on team satisfaction, project success, and contributor retention (Lim and Ling, 2002; Mcdonald et al., 2014). By analyzing real-world implementations of fairness frameworks, researchers could identify which approaches yield the most positive outcomes and how to adapt them to different team dynamics.
7.4. Satisfaction with Managers
@@
Previous studies support these findings. Suh & Hijal-Moghrabi (Suh and Hijal-Moghrabi, 2021) demonstrated that fair HR practices and fairness from superiors strongly impacted women’s job satisfaction. Choi & Rainey (Choi and Rainey, 2014) also found that diversity management and organizational fairness positively influenced job satisfaction, particularly for women, who reported higher satisfaction when fairness was perceived. Choi (Choi, 2017) further discovered that racial/ethnic minorities experienced higher job satisfaction when they held majority status in their organization, while Whites reported greater satisfaction in predominantly White settings. This highlights how ethnicity representation can impact one’s job satisfaction.
Gender and ethnic biases are prevalent SE as well. Gunawardena et al. (Gunawardena et al., 2022) showed that destructive criticism during code reviews negatively affects female developers’ moods and willingness to collaborate. Trinkenreich et al. (Trinkenreich et al., 2022b) found that women are sometimes assigned to projects based on gender rather than ability, while Paul et al. (Paul et al., 2019) revealed that male developers are more likely to write negative comments and offer less assistance to female colleagues. Our study builds on this body of work, which suggests that managers can play a crucial role in addressing gender bias. By prioritizing fair outcomes, clear communication, and unbiased decision-making processes, managers can significantly improve female developers’ satisfaction.
Similarly, developers from underrepresented ethnic groups, who are often a minority in their teams, face additional challenges in feeling heard and valued. Our findings show that for these developers, satisfaction with managers is strongly influenced by how inclusive and transparent managerial practices are, particularly in decision-making and providing opportunities to voice concerns or appeal decisions. When underrepresented developers feel they lack equal opportunities to participate, they may perceive this as unfair treatment, leading to dissatisfaction with their managers. This sense of disadvantage may stem from a lack of inclusion and transparency (Behfar et al., 2006), which can cause these developers to feel excluded from important team decisions.
This aligns with research by Hawes (Hawes and Wang, 2022), which found that minority workers’ job satisfaction improves significantly with strong supervisory support. In the predominantly White-dominated field of SE, Nadri et al. (Nadri et al., 2021) reported that less than 10% of contributions to open-source projects come from visibly non-White developers. These systemic barriers further emphasize the need for inclusive and transparent managerial practices to ensure underrepresented developers feel empowered and valued in their roles.
@@
7.5. Policy Barriers for Women and Minorities
@@
Policies that lack transparency or rely on subjective criteria can disproportionately disadvantage female and underrepresented developers. These groups often lack access to informal networks that help advance careers, facing a glass ceiling despite their qualifications (Daley, 1996). Even when women and minorities meet formal criteria like performance ratings, they often struggle to advance because they are excluded from the informal mentorship or sponsorship opportunities that typically benefit others, particularly white males (Daley, 1996). Unconscious biases continue to hinder their progress in Science, Technology, Engineering, and Mathematics (STEM) fields, limiting their opportunities for promotion and career growth, even in workplaces that have diversity policies in place (Greider et al., 2019; Villablanca et al., 2017).
Inflexible policies around working hours, parental leave, or caregiving responsibilities can also disproportionately impact female developers. Research from sectors like healthcare shows that work-life balance strongly affects organizational commitment and job satisfaction for women (Shabir and Gani, 2020). Similarly, remote work policies pose challenges if they are not equally accessible or if there is stigma attached to using them. Women, particularly those using flexible work options, often face a flexibility stigma, which can lead to perceptions of being less committed (Chung, 2018; Rogier and Padgett, 2004; Vandello et al., 2013). The COVID-19 pandemic has reduced some of the stigma around remote work, offering an opportunity to create more inclusive environments for women and minority employees (Barhate and Hirudayaraj, 2021).
@@
7.6. Managing Work Hours for the Less Experienced
@@
Several factors may contribute to the challenges faced by less experienced developers. One key challenge is the experience gap. Junior employees often struggle with unfamiliarity with company processes, technical skills, and organizational culture, which delays their performance and puts them at a disadvantage compared to senior employees who have already navigated these hurdles (Rastogi et al., 2015). Experienced developers tend to introduce fewer defects, even when working on similar tasks, and are generally more efficient in their performance (Tsunoda et al., 2017). Their experience and knowledge of the project help them handle complex tasks more efficiently, often completing them faster and with better results (Fronza and Vlasenko, 2011a). In contrast, less experienced developers spend more time searching for information and understanding code, leading to slower work processes (Ando et al., 2015). As less experienced developers gain more experience, their performance gradually become more like those of experienced developers (Fronza and Vlasenko, 2011b).
Research shows that long and irregular working hours negatively impact mental capacity and work quality for all developers, regardless of experience. Extended periods of continuous work reduce efficiency (Rodriguez et al., 2018). Meyer et al. suggest developing better tools to help developers plan their workday and boost personal productivity (Meyer et al., 2017). Additionally, Rastogi et al. highlight the value of mentorship or buddy systems for new hires, which can help them quickly understand company processes and become productive faster (Rastogi et al., 2015). A supportive environment where new employees feel comfortable asking questions without fear of judgment is also important for their development (Rastogi et al., 2015).
@@
7.7. Supporting Work-Limited Developers
Our analysis shows that developers with work limitations are often disadvantaged, especially when it comes to compensation in high-pressure situations. According to the National Health Interview Survey (NHIS), a work limitation is defined as a condition that limits a person’s ability to perform certain tasks or activities required for their job (Altman, 2014). For software engineers and developers, this could mean physical, cognitive, or psychological challenges that may affect their performance, such as chronic illnesses, disabilities, or other long-term health conditions.
A study examining work limitations due to chronic health conditions found that such limitations often led to reduced productivity and lower wages (Choe and Baldwin, 2017). However, workers whose jobs were better matched to their limitations experienced better outcomes, including higher wages and longer job tenure (Choe and Baldwin, 2017). Similarly, Siu et al. found that employees with chronic diseases valued fair treatment, clear communication, and flexible accommodations, all of which positively impacted their job satisfaction and work performance (Siu et al., 2013).
Providing accommodations and adjusting work conditions for employees with chronic physical conditions has been shown to improve their self-efficacy and reduce fatigue (Varekamp et al., 2011). Lerner et al. further confirmed that employees with chronic health conditions face challenges in managing time, physical demands, and interpersonal tasks, often resulting in productivity loss (Lerner et al., 2003). This aligns with the importance of non-monetary compensation, such as flexible work hours and support services, as essential factors for job satisfaction among these employees (Lerner et al., 2003).
Developers with disabilities face additional challenges in the workplace, including issues with project management practices, workplace dynamics, and the accessibility of development tools. For example, visually impaired developers often struggle with the accessibility of development environments and tools, which negatively impacts their productivity (Johnson et al., 2022; Huff et al., 2020). Programs like Catalisa, which train and hire people with disabilities as software developers, underscore the importance of tailored training to improve inclusion and accessibility within the industry (Cardoso-Pereira et al., 2023).
@@ 7.7. Supporting Work-Limited Developers
Future research could explore the types of accommodations that are most effective in improving job satisfaction for software developers with work limitations. This could include case studies of specific organizations that implement flexible working conditions (Conradie and Klerk, 2019; Mulay, 2017; Scholarios and Marks, 2004), and how different types of accommodations (Joyce et al., 2010) (e.g., flexible hours, ergonomic tools, or assistive technology) impact the developers.
8. Threats to Validity
Construct Validity. Since perceptions of fairness and job satisfaction are subjective, individual interpretations of survey questions may vary based on personal experiences or biases. To minimize this, we used well-established frameworks, including Colquitt’s fairness theory (Colquitt, 2001) and Storey et al.’s job satisfaction model (Storey et al., 2021). These scales have been rigorously tested and help ensure that the questions measure what they are intended to. We also ensured anonymity in the survey and conducted a pilot test with six respondents to refine any unclear questions and adjust wording based on feedback.
@@ 8. Threats to Validity
Internal Validity. While our study highlights a relationship between fairness perceptions and job satisfaction, we acknowledge that it does not conclusively establish causality. However, this limitation is mitigated by both our calculation of the effect size and the extensive body of organizational justice literature, consistent with prior research across various industries and contexts (Mossholder et al., 1998; Wesolowski and Mossholder, 1997; Masterson et al., 2000; Judge and Colquitt, 2004; Konovsky and Cropanzano, 1991; Moorman, 1991; Colquitt et al., 2001). Given this robust theoretical grounding and the significant effect sizes of minimum 50% effect observed in our study, we argue that the direction of the relationship identified is aligned with well-established patterns in organizational behavior.